Use Lagrange multipliers to find the points on a given curve that are nearest the origin. (You are not given the function f but it will be the distance formula between the point(x,y) and the point given.)

Need a worked example please
Answer #1

Let g(x, y, z) is a given curve and d = f(x, y, z) is the distance between any point (x, y, z) on the curve and the origin.

Then, according to the method of Langrange's multipliers, the distance d = f(x, y, z) is minimum when

\nabla f=\lambda\nabla g \text{ subject to } g=0

Example:

Let us find the point on the plane on the plane X-y+32 = 1 which is closet to the oregion. Distance of an arbitrary point (2,According to the method of langrangers multipliers, the absolute minimum of the distance function d²= f(x, y, z) will occur aбе - Вд - (1,-1,3) from o Vf c Ayg Э) (22, 2y, 27) - A(-1,3) 2 (23, 2y, 22) = (2,-2, 35) *, 2х - -) x, 2y =-2 ) y=- 27 - 3, 5Putting the values of 2, 4,2 in ③. 출 - (출) + 3(2) -) = 0 2) +2 +92 - 2 2. 02 ) 112 -2 O 코 2 || 류 키 출해 JI Hence, R 즐 =급 2 다를템